# Auditing and Generating Synthetic Data with Controllable Trust Trade-offs

### v3

@@ Acknowledgments
This trustworthiness index driven model selection is applicable to any generative modeling method, however for the sake of demonstrating a concrete and useful example, we focus on instrumenting transformers models with it. Owing to their versatility, transformer models and their variants provide state of the art generation quality across the different modalities we consider here, namely tabular, time series, and natural language. We refer to transformer models trained under trust constraints and selected via our trustworthiness index as TrustFormer. Figures 4(a) and 4(b) showcases model selection of TabFormer GPT models [3] ( private and non-private, TF(p,$\varepsilon=1$) and TF(n-p) respectively), based on the validation trustworthiness index corresponding to trade-off weights $\omega$. We see that indeed this trustworthiness index driven cross-validation leads to different model checkpoints being selected as $\omega$ varies, translating to an improved trustworthiness index when evaluated on real test data, thereby controlling the trust trade-offs of the resulting synthetic dataset. Note that this trustworthiness index selection is not only applicable for checkpoint selection, for example, given a particular checkpoint one can select the sampling strategy and other hyper-parameters of the inference that lead to the highest validation trustworthiness index.
@@
Appendix A Implementation Details for Training TrustFormers
A.1 Feature Extractor
Table 1 provides for each dataset the list of hyperparameters used in training for RoBERTa models used as an embedding E in the auditing.
@@ A.1 Feature Extractor
Table 1: Feature extractor RoBERTa models: columns are $N_{e}$ for number of epochs, $|V|$ for vocab size, $\lambda_{r}$ for learning rate, $b_{s}$ for batch size, $\mathrm{h_{F}}$ for field hidden size, $\mathrm{AH_{F}}$ for field number of attention heads, $h$ for hidden size, AH for number of attention heads, window size, stride, and $p_{\text{mlm}}$ for probability for masked language model. Please note that for MIMIC-III Notes dataset, our feature extractor is a pre-trained BioGPT-Large model.
A.2 GPT
Table 2 provides for each dataset the list of hyperparameters used in training for GPT models.
@@ A.2 GPT
MIMIC-III ${}^{\dagger}$ 57,716 $10^{-5}$ 16 10 no
Table 2: Generator GPT models: columns are $|V|$ for vocab size, $\lambda_{r}$ for learning rate, $b_{s}$ for batch size, window size, stride, $N_{e}$ for number of epochs, DP enabled (yes/no), DP $\epsilon$ budget, and $\max(||\Delta||)$ for DP max gradient norm.$\dagger$ for distributed training on 8 GPUs.
A.3 Details For Compute Environment and Distributed Training
@@
A.4 Timing of Training, Auditing and Aggregation
Table 3 provides timings for training both feature extractor and generator models, auditing, and aggregation of results for all our datasets. One can notice that Differential Privacy trainings comes at a computational cost making the training time per epoch longer.
@@ A.4 Timing of Training, Auditing and Aggregation
Table 3: Timings for trainings of our feature extractor models (Fea Ext), GPT models for both non-private and private trainings, audit of synthetic data, and aggregation of results. All training timings are given as hours per epoch while audit timings and aggregation are given in hours.
A.5 Private Sampling from TabGPT
@@
Appendix B Baselines Training Details
Multiple synthetic data generation techniques were employed to generate a varied set of baseline synthetic datasets to compare against TrustFormer generated synthetic datasets for quality assessment. Non private as well as diferentially private baselines were generated as seen in Table 3.
B.1 Non Private Baselines
@@ B.2 Private Baselines
Appendix C Implementation Details For Metrics
@@ C.4 Fairness
$\mathrm{TPR}_{D^{\prime}}(f)=\frac{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(f(\mathbf{x}^{(i)})=1)\mathbb{1}(y^{(i)}=1)}{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(y^{(i)}=1)}$ $\mathrm{FPR}_{D^{\prime}}(f)=\frac{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(f(\mathbf{x}^{(i)})=1)\mathbb{1}(y^{(i)}=0)}{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(y^{(i)}=0)}$
@@
$\Delta_{\mathrm{TPR}}=\mathrm{TPR}_{D_{privileged}}-\mathrm{TPR}_{D_{unprivileged}}$ $\Delta_{\mathrm{FPR}}=\mathrm{FPR}_{D_{privileged}}-\mathrm{FPR}_{D_{unprivileged}}$
@@ C.5 Robustness
Appendix D Implementation Details for Aggregation and trustworthiness index Scoring
@@
The empirical CDF (ECDF) estimation for each metric described in our auditing framework was implemented using Python statsmodels package222www.statsmodels.org. We used same set of ECDF mappings (created from same data pool) across all ranked items to ensure consistency. The $\beta$ weights in Eq. 7 were set to uniform values in our experiments, in absence of an a-priori preference over the individual metrics. Table 4 gives the total number ($M_{T}$) of aggregated metrics within each trust dimension.
In order to determine optimum configuration for a particular $\omega$ combination of TF trustworthiness index we performed steps as follows: (1) Using the development data partition, in each fold we evaluated trustworthiness indexes across all available epochs by pooling ECDF statistics across all such folds and epochs of the individual model at hand. (2) For each $\omega$ listed in Table 2, we stored the set of optimum epoch per fold. (3) We picked data from the test partition of the corresponding epoch of each fold. (3) A union set of test epochs from previous step was formed determining the final set of TF "models," some of which were optimum for multiple $\omega$ configurations (as can be seen in trustworthiness index tables, e.g., Table 9). Note that, after the final set of epochs and TF models was determined, the final rankings were based on ECDF estimated from pooled statistics of all TF configurations as well as baseline models.
@@
Table 4: Total metrics aggregated within each trust dimensions. <seed> refers to the 5 seeds considered in training MLP classifiers.
@@
Appendix E Independent Copula Versus Gaussian Copula
@@
Appendix F Trustworthiness Index Driven Cross Validation
(a) Controllable trust trade-offs via trustworthiness index driven cross-validation: Private TrustFormer TF(p,$\varepsilon=1$) (Bank Marketing dataset.)
(b) Controllable trust trade-offs via trustworthiness index driven cross-validation: Non-private TrustFormer TF(n-p) (Bank Marketing dataset.)
Figure 5: Controllable trust trade-offs. In the first row of Figures 4(a) (Private TrustFormer) and 4(b) (Non-Private TrustFormer), we depict the evolution of the validation trust dimensions indices (given in Equation 13) as function of the training time (epoch) within the training of TF on 5 different real data splits (80% train, 10% validation, 10% test, referred to in the plot as a trial). For the private case, we see that the fidelity index on average improves as training proceeds, while privacy index deteriorates on average as the training progresses. For the non-private case we see that while fidelity improves as training progresses, privacy has a sweet spot in the middle of the training. For both private and non private training, validation utility, fairness and robustness indices have different dynamics fluctuating around their geometric mean and have different peaks of performance within each trial. In the second row, we show for a trust trade-off weight ($\omega$) the validation trustworthiness index given in Equation (14), for each TF trained on a real data split as function of the training time (epoch). We plot in each panel, the validation trustworthiness index for $5$ different trade-offs weights $\omega_{all},\omega_{PUR}$,$\omega_{UFR}$, $\omega_{U},\omega_{e(PUF)}$ that are given in Table 2. We see that the selected epoch or checkpoint $t^{*}$ (given in Equation 15) is different for each training split and for each trade-off weight. This model selection, using the trustworthiness index allows the alignment of the generative model with the desired trust trade-off.
Appendix G Example of Auditing Report
@@ Appendix H Pseudo-code
Appendix I Additional Experimental Results
I.1 Bank Marketing Dataset
@@
Table 5: Bank Marketing: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
Table 6: Bank Marketing downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
Table 7: Bank Marketing dataset ranking. We see that overall TrustFormers that use trustworthiness index driven selection corresponding to the desired trade-offs outperform other synthetic data, across these desired trade-offs defined in Table 2. Note that the ranking here is on the mean trustworthiness index, where downstream tasks are evaluated on the test data in each real data fold.
@@
Table 8: Bank Marketing dataset. Ranking with uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (12) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index.
I.2 Recruitment Dataset
@@
Table 9: Recruitment Dataset: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
Table 10: Recruitment downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
Table 11: Recruitment dataset ranking. We see that overall TrustFormers that use trustworthiness index driven cross-validation corresponding to the desired trade-offs outperform other synthetic data, across these desired trade-offs defined in Table 2. Note that the ranking here is on the mean trustworthiness index, where downstream tasks are evaluated on the test data in each real data fold.
@@
Table 12: Recruitment dataset. Ranking under uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (12) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index.
I.3 Law School
@@
Table 13: Law school dataset: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
Table 14: Law School downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
Table 15: Law school dataset synthetic data ranking .We see that overall TrustFormers that use trustworthiness index driven cross-validation corresponding to the desired trade-offs outperform other synthetic data, on 6 out of 10 weighting in Table 2. Note that the ranking here is on the mean trustworthiness index , where downstream tasks are evaluated on the test data in each real data fold.
@@
Table 16: Law School dataset. Ranking under uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (12) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index. TrustFormer synthetic data is back to the top of the leader-board when we take into account uncertainty.

### v4

@@ Acknowledgments
This trustworthiness index driven model selection is applicable to any generative modeling method, however for the sake of demonstrating a concrete and useful example, we focus on instrumenting transformers models with it. Owing to their versatility, transformer models and their variants provide state of the art generation quality across the different modalities we consider here, namely tabular, time series, and natural language. We refer to transformer models trained under trust constraints and selected via our trustworthiness index as TrustFormer. Figures 7 and 7 showcases model selection of TabFormer GPT models [82] ( private and non-private, TF(p,$\varepsilon=1$) and TF(n-p) respectively), based on the validation trustworthiness index corresponding to trade-off weights $\omega$. We see that indeed this trustworthiness index driven cross-validation leads to different model checkpoints being selected as $\omega$ varies, translating to an improved trustworthiness index when evaluated on real test data, thereby controlling the trust trade-offs of the resulting synthetic dataset. Note that this trustworthiness index selection is not only applicable for checkpoint selection, for example, given a particular checkpoint one can select the sampling strategy and other hyper-parameters of the inference that lead to the highest validation trustworthiness index.
@@
-B Implementation Details for Training TrustFormers
-C Feature Extractor
Table I provides for each dataset the list of hyperparameters used in training for RoBERTa models used as an embedding E in the auditing.
@@
TABLE I: Feature extractor RoBERTa models: columns are $N_{e}$ for number of epochs, $|V|$ for vocab size, $\lambda_{r}$ for learning rate, $b_{s}$ for batch size, $\mathrm{h_{F}}$ for field hidden size, $\mathrm{AH_{F}}$ for field number of attention heads, $h$ for hidden size, AH for number of attention heads, window size, stride, and $p_{\text{mlm}}$ for probability for masked language model. Please note that for MIMIC-III Notes dataset, our feature extractor is a pre-trained BioGPT-Large model.
-D GPT
Table II provides for each dataset the list of hyperparameters used in training for GPT models.
@@
MIMIC-III Notes 57,716 $10^{-5}$ 16 10 no
TABLE II: Generator GPT models: columns are $|V|$ for vocab size, $\lambda_{r}$ for learning rate, $b_{s}$ for batch size, window size, stride, $N_{e}$ for number of epochs, DP enabled (yes/no), DP $\epsilon$ budget, and $\max(||\Delta||)$ for DP max gradient norm.$\dagger$ for distributed training on 8 GPUs.
-E Details For Compute Environment and Distributed Training
@@
-F Timing of Training, Auditing and Aggregation
Table III provides timings for training both feature extractor and generator models, auditing, and aggregation of results for all our datasets. One can notice that Differential Privacy trainings comes at a computational cost making the training time per epoch longer.
@@
TABLE III: Timings for trainings of our feature extractor models (Fea Ext), GPT models for both non-private and private trainings, audit of synthetic data, and aggregation of results. All training timings are given as hours per epoch while audit timings and aggregation are given in hours.
-G Private Sampling from TabGPT
@@
-H Baselines Training Details
Multiple synthetic data generation techniques were employed to generate a varied set of baseline synthetic datasets to compare against TrustFormer generated synthetic datasets for quality assessment. Non private as well as diferentially private baselines were generated as seen in Table IV.
-I Non Private Baselines
@@
-K Implementation Details For Metrics
@@
$\mathrm{TPR}_{D^{\prime}}(f)=\frac{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(f(\mathbf{x}^{(i)})=1)\mathbb{1}(y^{(i)}=1)}{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(y^{(i)}=1)}$
$\mathrm{FPR}_{D^{\prime}}(f)=\frac{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(f(\mathbf{x}^{(i)})=1)\mathbb{1}(y^{(i)}=0)}{\sum_{i=1}^{|D^{\prime}|}\mathbb{1}(y^{(i)}=0)}$
@@
$\Delta_{\mathrm{TPR}}$ $=\mathrm{TPR}_{D_{privileged}}-\mathrm{TPR}_{D_{unprivileged}}$ $\Delta_{\mathrm{FPR}}$
$=\mathrm{FPR}_{D_{privileged}}-\mathrm{FPR}_{D_{unprivileged}}$
@@
-Q Implementation Details for Aggregation and trustworthiness index Scoring
@@
The empirical CDF (ECDF) estimation for each metric described in our auditing framework was implemented using Python statsmodels package222www.statsmodels.org. We used same set of ECDF mappings (created from same data pool) across all ranked items to ensure consistency. The $\beta$ weights in Eq. 6 were set to uniform values in our experiments, in absence of an a-priori preference over the individual metrics. Table IV gives the total number ($M_{T}$) of aggregated metrics within each trust dimension.
In order to determine optimum configuration for a particular $\omega$ combination of TF trustworthiness index we performed steps as follows: (1) Using the development data partition, in each fold we evaluated trustworthiness indexes across all available epochs by pooling ECDF statistics across all such folds and epochs of the individual model at hand. (2) For each $\omega$ listed in Table II, we stored the set of optimum epoch per fold. (3) We picked data from the test partition of the corresponding epoch of each fold. (3) A union set of test epochs from previous step was formed determining the final set of TF "models," some of which were optimum for multiple $\omega$ configurations (as can be seen in trustworthiness index tables, e.g., Table IX). Note that, after the final set of epochs and TF models was determined, the final rankings were based on ECDF estimated from pooled statistics of all TF configurations as well as baseline models.
@@
TABLE IV: Total metrics aggregated within each trust dimensions. <seed> refers to the 5 seeds considered in training MLP classifiers.
@@
-R Independent Copula Versus Gaussian Copula
@@
-S Trustworthiness Index Driven Cross Validation
Figure 5: Controllable trust trade-offs via trustworthiness index driven cross-validation: Private TrustFormer TF(p,$\varepsilon=1$) (Bank Marketing dataset.)
Figure 6: Controllable trust trade-offs via trustworthiness index driven cross-validation: Non-private TrustFormer TF(n-p) (Bank Marketing dataset.)
Figure 7: Controllable trust trade-offs. In the first row of Figures 7 (Private TrustFormer) and 7 (Non-Private TrustFormer), we depict the evolution of the validation trust dimensions indices (given in Equation 12) as function of the training time (epoch) within the training of TF on 5 different real data splits (80% train, 10% validation, 10% test, referred to in the plot as a trial). For the private case, we see that the fidelity index on average improves as training proceeds, while privacy index deteriorates on average as the training progresses. For the non-private case we see that while fidelity improves as training progresses, privacy has a sweet spot in the middle of the training. For both private and non private training, validation utility, fairness and robustness indices have different dynamics fluctuating around their geometric mean and have different peaks of performance within each trial. In the second row, we show for a trust trade-off weight ($\omega$) the validation trustworthiness index given in Equation (13), for each TF trained on a real data split as function of the training time (epoch). We plot in each panel, the validation trustworthiness index for $5$ different trade-offs weights $\omega_{all},\omega_{PUR}$,$\omega_{UFR}$, $\omega_{U},\omega_{e(PUF)}$ that are given in Table II. We see that the selected epoch or checkpoint $t^{*}$ (given in Equation 14) is different for each training split and for each trade-off weight. This model selection, using the trustworthiness index allows the alignment of the generative model with the desired trust trade-off.
-T Example of Auditing Report
@@
-V Additional Experimental Results
-W Bank Marketing Dataset
@@
TABLE V: Bank Marketing: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
TABLE VI: Bank Marketing downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
TABLE VII: Bank Marketing dataset ranking. We see that overall TrustFormers that use trustworthiness index driven selection corresponding to the desired trade-offs outperform other synthetic data, across these desired trade-offs defined in Table II. Note that the ranking here is on the mean trustworthiness index, where downstream tasks are evaluated on the test data in each real data fold.
@@
TABLE VIII: Bank Marketing dataset. Ranking with uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (11) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index.
-X Recruitment Dataset
@@
TABLE IX: Recruitment Dataset: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
TABLE X: Recruitment downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
TABLE XI: Recruitment dataset ranking. We see that overall TrustFormers that use trustworthiness index driven cross-validation corresponding to the desired trade-offs outperform other synthetic data, across these desired trade-offs defined in Table II. Note that the ranking here is on the mean trustworthiness index, where downstream tasks are evaluated on the test data in each real data fold.
@@
TABLE XII: Recruitment dataset. Ranking under uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (11) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index.
-Y Law School
@@
TABLE XIII: Law school dataset: trust dimension indices. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
@@
TABLE XIV: Law School downstream task evaluation. Study on the effect of debiasing in utility training on trust indices of utility, fairness and robustness.
@@
TABLE XV: Law school dataset synthetic data ranking .We see that overall TrustFormers that use trustworthiness index driven cross-validation corresponding to the desired trade-offs outperform other synthetic data, on 6 out of 10 weighting in Table II. Note that the ranking here is on the mean trustworthiness index , where downstream tasks are evaluated on the test data in each real data fold.
@@
TABLE XVI: Law School dataset. Ranking under uncertainty of all considered synthetic data using $R^{\alpha}_{\tau}$ defined in Equation (11) for $\alpha=0.1$. We see that TrustFormer synthetic data is still aligned with the prescribed safeguards while having low volatility in its trustworthiness index. TrustFormer synthetic data is back to the top of the leader-board when we take into account uncertainty.